You are an assistant to the CFO and have collected the following data to conduct the analysis.

  • The company is subjected to a marginal tax rate of 35%.
  • The company can issue a 20-year, 7.6% semi-annual coupon bond at $1,219. New bonds would be privately placed with no floatation cost.
  • The company’s preferred stock currently sells for $30 per share. It pays a fixed dividend of $1.8 per share.
  • The company’s common stock currently sells for $40 per share. The most recently paid dividend was $1 per share. Dividends are expected to grow at a constant rate of 5% in the foreseeable future.
  • A flotation cost of 10% would be required to issue new common stock.
  • The company’s stock beta is 1.875, the market risk premium is 3% and the risk-free rate is 2%.
  • The company’s capital structure consists of 30% debt, 5% preferred stock, and 65% common equity.

(1)   Estimate the cost of the following capital components

  • Cost of Debt
  • Cost of Preferred Stock
  • Cost of Retained Earnings—based on DCF approach
  • Cost of Retained Earnings—based on CAPM approach
  • Cost of New Common Stock ---based on DCF approach

(2)   Calculate the firm’s WACC, assuming it does not want to issue new common stock.

(3)   Calculate the firm’s WACC, assuming the firm expands so rapidly that it must issue new common stock.

"not using excel"

Step by Step please.
Answer #1

a) Cost of debt can be calculated using I/Y function on a calculator

N = 20 x 2 = 40, PMT = 7.6% x 1000 / 2 = 38, PV = -1219, FV = 1000 => Compute I/Y = 2.87% (semi-annual)

Annual cost of debt = 2.87% x 2 = 5.74%

b) Cost of preferred = Dividend / Price = 1.8 / 30 = 6.00%

c) Using DCF,

Cost of retained earnings = D0 x (1 + g) / P + g = 1 x (1 + 5%) / 40 + 5% = 7.625%

d) Using CAPM,

Cost of retained earnings = Rf + beta x MRP = 2% + 1.875 x 3% = 7.625%

e) Cost of new stock = D0 x (1 + g) / (P x (1 - f)) + g = 1 x 1.05 / (40 x (1 - 10%)) + 5% = 7.92%

f) WACC = wd x rd x (1 - tax) + wp x rp + we x re

= 30% x 5.74% x (1 - 35%) + 5% x 6% + 65% x 7.625%

= 6.38%

g) WACC = 30% x 5.74% x (1 - 35%) + 5% x 6% + 65% x 7.92% = 6.57%